OpenStudy (anonymous):

Any Doctor Who fans out there will get this. I have 5 roses, then a girl named Rose(Tyler) gives me 5 more roses. How many roses do I have total? A. 9Rose B. 10Rose Whovians, see what I did there?

OpenStudy (hhopke):

Not a whovian (I know, sorry.) What is the answer?

OpenStudy (anonymous):

@hhopke, B: I will have 10 roses, 5 + 5 = 10 and 10Rose is my ship of DW. FORGET 9ROSE!!!!

OpenStudy (tkhunny):

If the extra 5 roses win your heart, doesn't the other Rose count? 11?

OpenStudy (anonymous):

No, a girl named Rose(Tyler) gave me 5 roses, and I already had the other 5, so 10

OpenStudy (tkhunny):

You didn't answer my objection.

OpenStudy (anonymous):

@tkhunny, because I don't understand your question.

OpenStudy (anonymous):

Or objection, sorry

OpenStudy (tkhunny):

You have 5 roses. Rose gives you 5 Roses. 5+5 = 10 You are so impressed with the five additional roses that you fall in love with Rose and live happily ever after. 10 + 1 = 11

OpenStudy (anonymous):

I like your judgement, but sadly no. After she gives you the roses, she gets trapped in another dimension and you never see her again. :( DOOMSDAY!!!!

OpenStudy (tkhunny):

Alas! Those dimensional portals can be like that. Too bad.

OpenStudy (anonymous):

I FREAKING HATE DALEKS!!!! Oh, and Cybermen, those too.